Solution and Answer please, Thank you!
- How long will it take to fill a reservoir with intake pipes A and B, if the reservoir can filled by A alone in 5 days and by B alone in 3 days.

Answer by richwmiller(17219)

1/a+1/b=1/x
1/3+1/5=1/x
5/15+3/15=1/x
8/15=1/x
x=15/8
x=1.875 days working together
check
1/a+1/b=1/x
1/3+1/5=1/x
5/15+3/15=1/x
8/15=1/x
15/8=x
x=1.875 days working together
ok